OpenStudy (anonymous):

Roberta is wrapping gift boxes. Each gift box is a rectangular prism with a square base of side 8 inches. The difference between the heights of the two gift boxes is 9 inches. If the larger box has a surface area of 544 square inches, what is the difference in the surface area of the two boxes? 576 square inches 648 square inches 324 square inches 288 square inches

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

answer is 288

OpenStudy (anonymous):

2*64+32h is the formula for prism. Bigger box:- 2*64+32h=544 32h=544-128 h=416/32 =13 smaller box:- 2*64+32(4)=256 Difference= 544-256=288